OpenStudy (anonymous):

Simplify the square root of 2 and 108

OpenStudy (anonymous):

sqrt(2) =1.41421 sqrt(108) = 10.39

OpenStudy (radar):

\[\sqrt{108}=\sqrt{36}\sqrt{3}=6\sqrt{3}\]

OpenStudy (radar):

\[\sqrt{2}=\sqrt{2}\] Can't be further simplied. It is an irrational value. It is an important value used a lot in electronics, converting ac voltage to its peak value comes to mind.